psirockin2:

Is trigonometry and is urgent I have to graph a harmonic function with S(0) equaling 2in. and the period equaling 0.5 seconds

Vocaloid:

I'm an idiot can you remind me what S(0) means in this context, is that the initial value

psirockin2:

I´m pretty sure S(0) is the equalibrium point?

psirockin2:

equilibrium... me no spell good

Vocaloid:



Vocaloid:

something like this? :S @sillybilly123 @Nnesha sorry to rely on you guys so much today, I'm just out of it

psirockin2:

I don´t know... I´m just as confused

psirockin2:

Wait... the formula! s(t)=acoswt w is omega a is the amp t is time...

psirockin2:

amp is the 2 inches... period is 0.5 sec so 2pi/omega has to equal 0.5 4pi? @Vocaloid Sorry to bug you

sillybilly123:

that's it Voca with the subsequently added data, we have \(s(t) = 2 ~ \cos \omega t \) and \(\omega = \dfrac{2 \pi}{0.5} = 4 \pi\)

sillybilly123: